OpenStudy (anonymous):

The length of the arc formed by the curve y = 7x^4 + 8x^8 between the points (3, 53055) and (8, 134246400) is represented by the integral where, a = ? and b = ? and f (x) = ?

OpenStudy (anonymous):

I like this question. :)

OpenStudy (anonymous):

It's tough I do not know where to start :(

OpenStudy (anonymous):

first you must get dy/dx

OpenStudy (anonymous):

Can you show me

OpenStudy (anonymous):

\[28x ^{3}+64x ^{7}\]

OpenStudy (anonymous):

derivative the given function with respect to x. You should arrive to that answer.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then here's the equation \[\int\limits_{3}^{8}\sqrt{1+(28x ^{3}+64x ^{7})^{2}}dx\]

OpenStudy (anonymous):

now what integrate

OpenStudy (anonymous):

can you do that step by step

OpenStudy (anonymous):

here comes the hard part. Can't we just arrive to the answer? :P Let me think.

OpenStudy (anonymous):

I figured it out man its a = 3 and b = 8

OpenStudy (anonymous):

f(x) = sqrt(1+ ( 28x^3+64x^7)^2)

OpenStudy (anonymous):

ya, you got it? cool. :)

OpenStudy (anonymous):

ok no problem